OpenStudy (anonymous):

how would you solve t ln(t)- t > 0?

OpenStudy (anonymous):

t>0,because t is in the In()

OpenStudy (anonymous):

start with \(t(\ln(t)-1)>0\)

OpenStudy (anonymous):

I simplified that to ln t>1...is that correct?

OpenStudy (anonymous):

that is not really "simpler" if you want to solve this inequality

OpenStudy (anonymous):

t > e^1

OpenStudy (anonymous):

you have \[t(\ln(t)-1)>0\] and we know that the domain of the log is \(t>0\) so the first factor is always positive then it comes down to solving \[\ln(t)-1>0\implies \ln(t)>1\implies t>e\]

OpenStudy (anonymous):

oh i am sorry i did not read your answer carefully yes, it comes down to \(\ln(t)>1\) as you said